• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 15:01:47 on April11,2007

    Entry number 198635

    keyword=target adjusted from 7.5 to 8.0 K

    Following instructions from RC / JP, I have changed the pid set point so the target operates at 8.0 K. I also increased the hi and hihi alarm points by 0.5 K. The pressure is pretty stable at ~197 psi. The heat load with beam now off is ~450 W from the high power heater.

    FIGURE 1